OpenStudy (anonymous):

I have a 95 grade point average. I have submitted an assignment that is 31 % of my grade. What would be the maximum amount of points that could be deducted from my grade point average?

OpenStudy (ybarrap):

New GPA (g): $$ \large{ g_{new}=\Sigma \cfrac{g_i+0.31}{n+1}\\ g_{old}=\Sigma \cfrac{g_i}{n}\\ g_{new}-g_{old}=\Sigma \cfrac{g_i+0.31}{n+1}-\Sigma \cfrac{g_i}{n}\\ =\Sigma \cfrac{(g_i+0.31)n-g_i(n+1)}{(n+1)n}\\ =\Sigma\cfrac{-g_i+0.31n}{n^2+n}\\ =\Sigma\cfrac{-g_i}{n^2+n}+\cfrac{0.31n}{n^2+n}\\ g_{old}-g_{new}=\cfrac{-0.31n}{n^2+n}+\Sigma\cfrac{g_i}{n^2+n}\\ =\cfrac{-0.31}{n+1}+\Sigma\cfrac{g_i}{n^2+n}\\ =\cfrac{-0.31}{n+1}+\cfrac{0.95}{n+1}\\ =\cfrac{0.95-0.31}{n+1}\\ =\cfrac{0.64}{n+1} } $$ Where \(g_i\) are the grades for each prior assignment and n are the number of prior assignments. Now, \(g_{old}-g_{new}\) is maximum n=1, since n=0 is not possible. $$ \large{ max(g_{old}-g_{new})=\cfrac{64}{2}=32 } $$ This is also the maximum number of points tat could be deducted from your GPA.